OpenStudy (anonymous):

sole the equation log(5)^x=6

OpenStudy (anonymous):

... apply laws of logs log x^y = y log x...

OpenStudy (anonymous):

If we have \[ (\log(5))^x=6 \]Then: \[ \log((\log(5))^x)=\log(6) \]That is: \[ x\log(\log(5))=\log(6) \]Which gives us: \[ x=\frac{\log(6)}{\log(\log(5))} \]